OpenStudy (anonymous):

sin 30 degree + cos 45 degree= (Simplify answer, type answer, using radicals as needed. Rationalize all denominators. Use integers of fractions for any numbers in the expression)

OpenStudy (anonymous):

sin (30) = 0.5 cos (45) = \[\frac{ 1 }{ \sqrt{2} } = \frac{ \sqrt{2} }{ 2 }\] These are things you'll just need to memorize. Cos (30) as well. Sqrt of 3 over 2

OpenStudy (anonymous):

you need to memorize the unit circle

OpenStudy (anonymous):

Is the result to this -0.4627 degree

OpenStudy (anonymous):